# PixT3: Pixel-based Table-To-Text Generation

### v2

@@ Abstract
Table-to-text generation involves generating appropriate textual descriptions given structured tabular data. It has attracted increasing attention in recent years thanks to the popularity of neural network models and the availability of large-scale datasets. A common feature across existing methods is their treatment of the input as a string, i.e., by employing linearization techniques that do not always preserve information in the table, are verbose, and lack space efficiency. We propose to rethink data-to-text generation as a visual recognition task, removing the need for rendering the input in a string format. We present PixT3, a multimodal table-to-text model that overcomes the challenges of linearization and input size limitations encountered by existing models. PixT3 is trained with a new self-supervised learning objective to reinforce table structure awareness and is applicable to open-ended and controlled generation settings. Experiments on the ToTTo Parikh et al. (2020a) and Logic2Text Chen et al. (2020c) benchmarks show that PixT3 is competitive and, in some settings, superior to generators that operate solely on text. [^1] Our code, models, and data are available at xxx.yyy.zzz.
@@ 1 Introduction
Problematically, representing all the table information in a linear sequence results in a verbose representation that often exceeds the context window limit of popular Transformer models Vaswani et al. (2017). The challenge of processing such long sequences has fostered the development of even more controlled methods which refrain from encoding the table as a whole, concentrating exclusively on highlighted content (e.g., only the yellow cells in Figure 1). Unfortunately, models trained on abridged input have difficulty generalizing to new domains while being practically ineffective in scenarios where content selection is not provided.
@@ 2 Related Work
Previous attempts to address table-to-text generation from a visual recognition perspective Dash et al. (2023); Srihari et al. (2003) have relied on OCR methods which first extract text from the image and then feed it as string to a generation model. Aside from being noisy, these techniques typically embrace a text-centric point of view, treating the image as a limitation rather than an informative modality. Our work builds on recent visual language understanding models Kim et al. (2022); Davis et al. (2022); Lee et al. (2023) which are based exclusively on pixels and have managed to outperform OCR methods in several natural language processing tasks Rust et al. (2023); Huang et al. (2022); Masry et al. (2022); Salesky et al. (2023).
A few efforts have recently explored multimodal approaches to processing tables for various tasks, including table-to-text generation. Dash et al. (2023) convert images into HTML tokens which are subsequently linearized and processed by a traditional text-to-text model. Other work Chen et al. (2023a) focuses on recognizing the structure of tables from images as an independent task. It also leverages multimodal pretraining and unsupervised table structure learning objectives, but ignores the content of table cells and their relations. To the best of our knowledge, our work is the first to conceptualize data-to-text generation as a visually-situated language understanding problem.
@@ 3 Problem Formulation
The task of table-to-text generation is to take a structured table $\bm{t}$ as input and output a natural language description $\bm{y}=[y_{1},\dots,y_{k}]$ where $k$ is the length of the description. Table $\bm{t}$ is typically reformatted as a sequence of textual records $\bm{t}=[t_{1,1},t_{1,2},\dots,t_{i,j},\dots,t_{m,n}]$ where $m$ and $n$ respectively denote the number of rows and columns of $\bm{t}$.
We approach the task from a visual recognition perspective, and expect the input table to be an image $\bm{x}$. The image is reshaped into a sequence of patches analogous to linguistic tokens. More formally, for an input image $\bm{x}\in R^{H\times W\times C}$ and patch size $p$, we create $N$ image patches denoted as $x_{p}\in R^{N\times(P^{2}\cdot C)}$. $(H,W)$ is the resolution of the original image, $C$ is the number of channels, $(P,P)$ is the resolution of each image patch, and $N=\frac{HW}{P^{2}}$ the resulting number of patches, which serves effectively as the input sequence length. Our proposed model learns to autoregressively estimate the conditional probability of a text sequence from a source image as:
@@ 4 The PixT3 Model
We initialize PixT3’s model weights with Pix2Struct; we next adopt a curriculum training strategy which instills in our model knowledge about tables and their structure (see Section 4.2); and finally, we fine-tune on table-to-text generation datasets such as ToTTo Parikh et al. (2020a) with a task-specific self-supervised objective.
@@ 4.2 Structure Learning Curriculum
Intermediate Training Dataset
@@
We determine the structure of each table (size, column, and row spans) randomly, following ToTTo’s training set distribution. We cap the generation process at a maximum of 20 columns and 75 rows. Table cells are filled with synthetic values consisting of a random combination of one to five random English alphabet characters and digits, functioning as identifiers rather than meaningful values (see Figure 3 for an example). Our intermediate training dataset contains 135,400 synthetic tables, 120,000 for training, 7,700 for validation, and 7,700 for testing.
@@
While masking is a widely adopted learning objective Devlin et al. (2019), it does not transfer easily to our table-to-text generation task; table values are not naturally correlated to neighboring values and thus a masked cell cannot be easily predicted from other cells in its context. Table values could be rearranged so that they correlate to their neighbors, however, early experiments showed that this type of objective does not improve downstream task performance (see Appendix D for details). Another common pretraining objective is table linearization Chen et al. (2023a), which, however, scales poorly with table size, leading to slow pretraining.
@@ 5 Experimental Setup
PixT3 models were fine-tuned with a batch size of 8 and a gradient accumulation of 32 for 14K steps on an A100 GPU. Checkpoints were selected according to best performance on the validation set. All models used an input sequence length of 2,048 patches and were optimized with AdamW Loshchilov and Hutter (2017). We used a learning rate scheduler with a linear warmup of 1,000 steps to 0.0001, followed by cosine decay to 0. The decoder maximum sequence length was set to 50 tokens, which covers 97.49% of the target descriptions in the training data. PixT3 was trained for 1.4k steps with the self-supervised objective described in Section 4.2. Our decoder was not frozen during this intermediate training, as initial experiments showed that a fully trained model outperformed one with frozen decoder weights.
@@
We evaluated PixT3 against several text-only models with similar parameter sizes. These include CoNT An et al. (2022), the top performant (published) model in the ToTTo leaderboard. [^3] A model named SKY appears to slightly outperform CoNT in the leaderboard, however, at the time of writing, we were not able to verify this, i.e., by finding a publication or preprint describing this model. CoNT is a text-to-text generation model which makes use of contrastive learning, through improved selection of contrastive examples, a new contrastive loss, and a global decoding strategy. CoNT expects the input table to be converted to a string, and is built on top of T5-base (220M parameters). We also compared against Lattice Wang et al. (2022), a model which enforces awareness of table layout though pruning the attention flow and encoding cells in a way that is invariant to their relative position in a sequence. Their model also uses T5-base and expects linearized input. In addition, we report results with vanilla T5-base which performed competitively on the ToTTo leaderboard without any task specific modifications Kale and Rastogi (2020); An et al. (2022). All comparison models and PixT3, were trained on the ToTTo training set in our three generation settings. [^4] Comparison models were trained with the authors’ publicly available scripts.
For our out-of-domain experiments, we also compare against LLaVA-1.5 Liu et al. (2023), a large pretrained multimodal model (13B parameters) which is built on top of the CLIP encoder Radford et al. (2021) and the LLaMA language decoder Touvron et al. (2023), and fine-tuned on vision-language instructions. LLaVa has not been fine-tuned specifically for table-to-text generation, however, it is interesting to see if sufficiently large scale is all it takes to do well on the table generation task. To obtain table descriptions, LLaVA was prompted with an image, an instruction, and three table descriptions for each generation setting (see Appendix F for details). We summarize the number of parameters for all comparison models in Table 2.
@@ 6 Results
Table 1 summarizes our results on ToTTo in our three generation settings. We evaluated model performance automatically with the same metrics used to rank participant systems in the ToTTo leaderboard. These include BLEU Papineni et al. (2002) which is as a proxy for fluency, PARENT Dhingra et al. (2019), a metric proposed specifically for data-to-text evaluation that takes the table into account, serving as a proxy of faithfulness, and BLEURT Sellam et al. (2020); the latter is a composite metric that takes a reference and model output as input, and returns a score that indicates the extent to which the output is fluent and conveys the meaning of the reference. Note that ToTTo features two splits in the development/test set containing tables whose header values are present (overlapping split) and absent (non-overlapping split) in the training. Results on the test set, which is not publicly available, were obtained via submitting to the ToTTo leaderboard.
@@
In Table 3 we perform an ablation study comparing PixT3 with and without our structure learning curriculum and self-supervised objective (Section 4.2). For both models we follow the same fine-tuning process: we render tables into images, identify the optimal point of image compression and truncation (see Section 4.1), and perform hyper-parameter search to optimize Pix2Struct-base for our task. Vanilla PixT3 (second row in Table 3) shows a substantial improvement over an out-of-the-box Pix2Struct model which achieves a BLEU score of 0.2 and PARENT score of 0.6 on the ToTTo development set. Adding the intermediate training curriculum (second row in Table 3) slightly improves vanilla PixT3 across evaluation metrics. Inspection of the descriptions produced by the two models revealed that they tend to generate descriptions that are semantically equivalent to the target (43% of the time). However, vanilla PixT3 is less faithful, especially in the OpenE setting, where it tends to disregard or misinterpret the structure of the table, e.g., by associating unrelated cells.
@@ 8 Limitations
While PixT3 shows promising results, its performance is affected by the dimension of the input tables (for instance, 16% of the Wikipedia tables in ToTTo remain too big for PixT3 to represent effectively). It would be interesting to look into alternative ways of preprocessing very large tables, e.g., by rendering them via multiple images. While our proposed intermediate training methodology mitigates faithfulness errors, the model still struggles with hallucinations, falling short of human-level performance. Finally, PixT3 is not able to perform reasoning, i.e., to infer information which is is not explicitly stated in the table or make logical connections between concepts. Aside from new training objectives, a promising direction would be to combine the visual representations with an intermediate planning component which encourages the model to reason about the input while generating the output.
@@ Appendix D Intermediate Training
T5-base -0.298 -0.395 -0.191 -0.293
T5-3B -0.309 -0.416 -0.194 -0.305
Lattice -0.287 -0.382 -0.195 -0.288
CoNT -0.293 -0.387 -0.190 -0.289
@@
T5-base -0.371 -0.458 -0.278 -0.368
T5-3B -0.385 -0.456 -0.301 -0.378
Lattice -0.377 -0.451 -0.302 -0.377
CoNT -0.370 -0.452 -0.281 -0.366
OpenE PixT3 -0.332 -0.414 -0.258 -0.336
@@ Appendix E Additional Results and Examples
LLaVA -1.230
T5-base -1.086
T5-3B -1.079
Lattice -1.060
CoNT -1.103
TControl PixT3 -1.104
LLaVA -1.189
T5-base -1.147
T5-3B -1.167
Lattice -1.147
CoNT -1.159
LControl PixT3 -1.073
LLaVA -1.184
T5-base -1.237
T5-3B -1.196
Lattice -1.231
CoNT -1.231
OpenE PixT3 -1.213

### v3

@@ Abstract
Table-to-text generation involves generating appropriate textual descriptions given structured tabular data. It has attracted increasing attention in recent years thanks to the popularity of neural network models and the availability of large-scale datasets. A common feature across existing methods is their treatment of the input as a string, i.e., by employing linearization techniques that do not always preserve information in the table, are verbose, and lack space efficiency. We propose to rethink data-to-text generation as a visual recognition task, removing the need for rendering the input in a string format. We present PixT3, a multimodal table-to-text model that overcomes the challenges of linearization and input size limitations encountered by existing models. PixT3 is trained with a new self-supervised learning objective to reinforce table structure awareness and is applicable to open-ended and controlled generation settings. Experiments on the ToTTo Parikh et al. (2020a) and Logic2Text Chen et al. (2020c) benchmarks show that PixT3 is competitive and, in some settings, superior to generators that operate solely on text. [^1] Our code, models, and data are available at https://github.com/alonsoapp/PixT3.
@@ 1 Introduction
Problematically, representing tabular information as a linear sequence results in a verbose representation that often exceeds the context window limit of popular Transformer models Vaswani et al. (2017). The challenge of processing such long sequences has fostered the development of even more controlled methods which refrain from encoding the table as a whole, concentrating exclusively on highlighted content (e.g., only the yellow cells in Figure 1). Unfortunately, models trained on abridged input have difficulty generalizing to new domains while being practically ineffective in scenarios where content selection is not provided.
@@ 2 Related Work
Previous attempts to address table-to-text generation from a visual recognition perspective Dash et al. (2023); Srihari et al. (2003) have relied on OCR methods which first extract text from the image and then feed it as a string to a generation model. Aside from being noisy, these techniques typically embrace a text-centric point of view, treating the image as a limitation rather than an informative modality. Our work builds on recent visual language understanding models Kim et al. (2022); Davis et al. (2022); Lee et al. (2023) which are based exclusively on pixels and have managed to outperform OCR methods in several natural language processing tasks Rust et al. (2023); Huang et al. (2022); Masry et al. (2022); Salesky et al. (2023).
The field of Vision Language Models (VLMs) has also experienced significant growth in recent years Liu et al. (2023); Ye et al. (2023b); Bai et al. (2023); Wang et al. (2023); Alayrac et al. (2022). While most of them focus primarily on natural images, a few are starting to explore the application of dual encoder architectures to visually represented language Ye et al. (2023a); Zhang et al. (2023). However, these architectures are not parameter lean (with increased model size of a factor of 40 or more compared to Pix2Struct), and some continue to rely on fixed resolution images which can be particularly problematic when processing tabular data.
A few other efforts have recently explored multimodal approaches to processing tables for various tasks, including table-to-text generation. Dash et al. (2023) convert images into HTML tokens which are subsequently linearized and processed by a traditional text-to-text model. Other work Chen et al. (2023a) focuses on recognizing the structure of tables from images as an independent task. It also leverages multimodal pretraining and unsupervised table structure learning objectives, but ignores the content of table cells and their relations. To the best of our knowledge, our work is the first to conceptualize data-to-text generation as a visually-situated language understanding problem.
@@ 3 Problem Formulation
The task of table-to-text generation aims to take a structured table $\bm{t}$ as input and output a natural language description $\bm{y}=[y_{1},\dots,y_{k}]$ where $k$ is the length of the description. Table $\bm{t}$ is typically reformatted as a sequence of textual records $\bm{t}=[t_{1,1},t_{1,2},\dots,t_{i,j},\dots,t_{m,n}]$ where $m$ and $n$ respectively denote the number of rows and columns of $\bm{t}$.
We approach this task from a visual recognition perspective, and expect the input table to be an image $\bm{x}$. The image is reshaped into a sequence of patches analogous to linguistic tokens. More formally, for an input image $\bm{x}\in R^{H\times W\times C}$ and patch size $p$, we create $N$ image patches denoted as $x_{p}\in R^{N\times(P^{2}\cdot C)}$. $(H,W)$ is the resolution of the original image, $C$ is the number of channels, $(P,P)$ is the resolution of each image patch, and $N=\frac{HW}{P^{2}}$ the resulting number of patches, which serves effectively as the input sequence length. Our proposed model learns to autoregressively estimate the conditional probability of a text sequence from a source image as:
@@ 4 The PixT3 Model
We initialize PixT3’s model weights with Pix2Struct; we next adopt a curriculum training strategy which instills in our model knowledge about tables and their structure (see Section 4.2); and finally, we fine-tune on table-to-text generation datasets such as ToTTo Parikh et al. (2020a) with a task-specific supervised objective.
@@ 4.2 Structure Learning Curriculum
Dataset for Intermediate Training
@@
We determine the structure of each table (size, column, and row spans) randomly, following ToTTo’s training set distribution. We cap the generation process at a maximum of 20 columns and 75 rows. Table cells are filled with synthetic values consisting of a random combination of one to five random English alphabet characters and digits, functioning as identifiers rather than meaningful values (see Figure 3 for an example). Our dataset contains 135,400 synthetic tables, 120,000 for training, 7,700 for validation, and 7,700 for testing.
@@
While masking is a widely adopted learning objective Devlin et al. (2019), it does not naturally transfer to our table-to-text generation task; table values are not naturally correlated to neighboring values and thus a masked cell cannot be easily predicted from other cells in its context. Table values could be rearranged so that they correlate to their neighbors, however, early experiments showed that this type of objective does not improve downstream task performance (see Appendix D for details). Another common pretraining objective is table linearization Chen et al. (2023a), which, however, scales poorly with table size, leading to slow pretraining.
@@
4.3 PixT3 Fine-tuning
The intermediately pre-trained PixT3 is subsequently fine-tuned on an image-rendered dataset (see Section 4.1). In experiments, we use ToTTo Parikh et al. (2020b), however, our approach is not tied to a particular style of tables. Due to our model’s requirement for unimodal input, we treat table-related information (such as its title) as part of the table itself and render them both as one image (see Lee et al. 2023 for a similar approach).
@@ 5 Experimental Setup
PixT3 models were fine-tuned with a batch size of 8 and a gradient accumulation of 32 steps on a single NVIDIA A100 80GB GPU. Checkpoints were selected according to best performance on the validation set. All models used an input sequence length of 2,048 patches and were optimized with AdamW Loshchilov and Hutter (2017). We used a learning rate scheduler with a linear warmup of 1,000 steps to 0.0001, followed by cosine decay to 0. The decoder maximum sequence length was set to 50 tokens, which covers 97.49% of the target descriptions in the training data. PixT3 was trained for 1.4k steps with the self-supervised objective described in Section 4.2. Our decoder was not frozen during intermediate training, as initial experiments showed that a fully trained model outperformed one with frozen decoder weights. A full list of fine-tuning hyper-parameters can be found in Appendix H.
@@
We evaluated PixT3 against several text-only models with similar parameter sizes. These include CoNT An et al. (2022), the top performant (published) model in the ToTTo leaderboard. [^3] A model named SKY appears to slightly outperform CoNT in the leaderboard, however, at the time of writing, we were not able to verify this, i.e., by finding a publication or preprint describing this model. CoNT is a text-to-text generation model which makes use of contrastive learning, through improved selection of contrastive examples, a new contrastive loss, and a global decoding strategy. CoNT expects the input table to be converted to a string, and is built on top of T5-base (220M parameters). We also compared against Lattice Wang et al. (2022), a model which enforces awareness of table layout though pruning the attention flow and encoding cells in a way that is invariant to their relative position in a sequence. This model also uses T5-base and expects linearized input. In addition, we report results with vanilla T5-base which performed competitively on the ToTTo leaderboard without any task specific modifications Kale and Rastogi (2020); An et al. (2022). All comparison models and PixT3, were trained on the ToTTo training set in our three generation settings. [^4] Comparison models were trained with the authors’ publicly available scripts.
For our out-of-domain experiments, we also compare against LLaVA-1.5 Liu et al. (2023), a large pretrained multimodal model (13B parameters) which is built on top of the CLIP visual encoder Radford et al. (2021) and the Vicuna-7B language model Zheng et al. (2023), and fine-tuned on vision-language instructions. LLaVA has not been fine-tuned specifically for table-to-text generation, however, it is interesting to see if sufficiently large scale is all it takes to do well on the table-to-text generation task. LLaVA can only handle a single image at each forward pass. This limitation prevents it from performing inference in an in-context learning setting, where the model has access to multiple input-output examples at the same time. To approximate in-context learning as closely as possible, we provided LLaVA with an image, an instruction, and three table descriptions as output examples for each generation setting (see Appendix F for details). We summarize the number of parameters for all comparison models in Table 2.
we do still provide a few description examples in our prompt to ensure a fair zero-shot comparison. All prompts used for LLaVA in this evaluation can be found in Appendix F.
@@ 6 Results
Table 1 summarizes our results on ToTTo in our three generation settings. We evaluated model performance automatically with the same metrics used to rank participant systems in the ToTTo leaderboard. These include BLEU Papineni et al. (2002) which is as a proxy for fluency, PARENT Dhingra et al. (2019), a metric proposed specifically for data-to-text evaluation that takes the table into account, serving as a proxy of faithfulness, and BLEURT Sellam et al. (2020); the latter is a composite metric that takes a reference and model output as input, and returns a score that indicates the extent to which the output is fluent and conveys the meaning of the reference. Note that ToTTo features two splits in the development/test set containing tables whose header values are present (overlapping split) and absent (non-overlapping split) in the training set. Results on the test set, which is not publicly available, were obtained via submitting to the ToTTo leaderboard.
@@
In Table 3 we perform an ablation study comparing PixT3 with and without our structure learning curriculum and self-supervised objective (Section 4.2). For both models we follow the same fine-tuning process: we render tables into images, identify the optimal point of image compression and truncation (see Section 4.1), and perform hyper-parameter search to optimize Pix2Struct-base for our task. Vanilla PixT3 (second row in Table 3) shows a substantial improvement over an out-of-the-box Pix2Struct model which achieves a BLEU score of 0.2 and PARENT score of 0.6 on the ToTTo development set. Adding the intermediate training curriculum (second row in Table 3) slightly improves vanilla PixT3 across evaluation metrics.
Manual inspection of the descriptions produced by the two PixT3 model variants reveals they are often semantically equivalent to the target (43% of the time). Nevertheless, the intermediate training curriculum substantially reduces structure-based faithfulness errors, especially in the OpenE setting. On a sample of 200 outputs (randomly selected from the development set), we found that 23% of the descriptions produced by vanilla PixT3 disregard or misinterpret the structure of the table. Structural faithfulness errors reduce to 7% when PixT3 is trained with our structure learning curriculum.
@@ 8 Limitations
While PixT3 shows promising results, its performance is affected by the dimension of the input tables (for instance, 16% of the Wikipedia tables in ToTTo remain too big for PixT3 to represent effectively). It would be interesting to look into alternative ways of preprocessing very large tables, e.g., by rendering them via multiple images. While our proposed intermediate training methodology mitigates faithfulness errors, the model still struggles with hallucinations, falling short of human-level performance.
Finally, PixT3, as well as other comparison systems, have limited reasoning capabilities, e.g., they cannot infer information which is not explicitly stated in the table or make logical connections between concepts. PixT3’s superior performance in terms of faithfulness on Logic2Text (see Table 4) is due to generating simpler sentences rather than superior reasoning skills. Thus, aside from new training objectives, a promising direction would be to combine the visual representations with an intermediate planning component that encourages the model to reason about the input while generating the output.
Acknowledgements
We thank the meta-reviewer and anonymous reviewers for their constructive feedback. The authors also thank Ander Salaberria for his insightful comments on earlier versions of this work. We gratefully acknowledge the support of the UK Engineering and Physical Sciences Research Council (grant EP/L016427/1), the Basque Government (Research group funding IT-1805-22), MCIN/AEI/10.13039/501100011033 project AWARE (TED2021-131617B-I00), European Union NextGenerationEU/PRTR, and the LUMINOUS project (HORIZON-CL4-2023-HUMAN-01-21-101135724).
@@ Appendix D Intermediate Training
T5-base --0.298 --0.395 --0.191 -0.293
T5-3B --0.309 --0.416 --0.194 -0.305
Lattice --0.287 --0.382 --0.195 -0.288
CoNT --0.293 --0.387 --0.190 -0.289
@@
T5-base --0.371 --0.458 --0.278 -0.368
T5-3B --0.385 --0.456 --0.301 -0.378
Lattice --0.377 --0.451 --0.302 -0.377
CoNT --0.370 --0.452 --0.281 -0.366
OpenE PixT3 --0.332 --0.414 --0.258 --0.336
@@ Appendix E Additional Results and Examples
LLaVA --1.230
T5-base --1.086
T5-3B --1.079
Lattice --1.060
CoNT --1.103
TControl PixT3 --1.104
LLaVA --1.189
T5-base --1.147
T5-3B --1.167
Lattice --1.147
CoNT --1.159
LControl PixT3 --1.073
LLaVA --1.184
T5-base --1.237
T5-3B --1.196
Lattice --1.231
CoNT --1.231
OpenE PixT3 --1.213
@@ Appendix G Human Evaluation Guidelines
Appendix H PixT3 Fine-tuning Hyper-parameters
PixT3 models across all three settings (TControl, LControl, OpenE) were fine-tuned using the same hyper-parameters. To prevent over-fitting, we employed early stopping based on the BLEU score computed on the validation set every 250 steps. Table 8 enumerates the specific hyper-parameter values used in PixT3, with all remaining parameters set to the default values defined in Pix2Struct Lee et al. (2023).
Hyperparameter Value
Optimizer AdamW
Learning rate 0.0001
Warm-up steps 1000
Max. input patches 2048
Shuffle train data False
Epochs 30
Train batch size 8
Gradient accum. steps 32
Mixed precision fp16
Evaluation batch size 32
Eval freq. steps 250
Inf. beam search 8 beams
Table 8: Hyperparameters used in PixT3.